OpenStudy (anonymous):

3w+6(w+1)-10=11w-3(4-2w) solve for w please, problem is pissing me off i keep getting -2...

OpenStudy (akshay_budhkar):

its w=1..

OpenStudy (akshay_budhkar):

3w+6w+6-10=11w-12+6w 8w=8 w=1

OpenStudy (mimi_x3):

3w +6w + 6 -10 = 11w-12+6w 9w -4 = 17w -12 -8w = 8 w= -1 ( not sure if its right though)

OpenStudy (anonymous):

it's 1 thanks though ;D

OpenStudy (anonymous):

Thanks ashkay!

OpenStudy (mimi_x3):

wtf i keep on getting -1 argghh